OpenStudy (anonymous):

1. Which of the following is equivalent to x^2 y^2/xy^2 when y=/O and x=/o ? (1 point) x^6y5^ x^5y x^4y x^4

OpenStudy (anonymous):

what is /O and /o?

OpenStudy (anonymous):

wait i will show you

OpenStudy (anonymous):



OpenStudy (anonymous):

did you get it

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

\[\frac{x^2y^2}{xy^2}\]...this is the problem?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Are they saying y=x?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

\[\frac{x^2y^2}{xy^2}=x\]

OpenStudy (anonymous):

I'm not following what you're trying to do with that symbol

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

can't give an answer other than that because i'm not following what that is

OpenStudy (anonymous):

how about when y=0 and x=0

OpenStudy (anonymous):

then it's 0

OpenStudy (anonymous):

it was the last one

OpenStudy (anonymous):

i guess it means x is not equal to zero and y is not equal to zero. in that case your answer will be... x